Last fall, a gardener planted 74 tulip bulbs. She found that only 62 of the bulbs bloomed in the spring.
a) Find the empirical probability that a tulip bulb of this type will bloom
b) How many of the bulbs should she plant next fall if she would like at least 95 to bloom?

Answer by stanbon(57226)

Last fall, a gardener planted 74 tulip bulbs. She found that only 62 of the bulbs bloomed in the spring.
a) Find the empirical probability that a tulip bulb of this type will bloom
Ans: 62/74= 0.837838...
====================================
b) How many of the bulbs should she plant next fall if she would like at least 95 to bloom?
---
Use a proportion:
x/95 = 74/62
x = (74/62)*95
x = 113.387097
Rounding up she should plant 114 plants
================================================
Cheers,
Stan H.